The Boxwood Company sells blankets for $60 each. The following was taken from the inventory records during May. The company had no beginning inventory on May 1.


Date Blankets Units Cost
May 3   Purchase 5 $20          
10   Sale 3
17   Purchase 10 $24          
20   Sale 6
23   Sale 3
30   Purchase 10 $30          



Assuming that the company uses the perpetual inventory system, determine the gross profit for the month of May using the LIFO cost method.

$452

$444

$356

$348
Answer #1

Cost of goods sold under LIFO = (3*20+9*24) = $276

Sales revenue = 12*60 = 720

Gross profit = Sales-Cost of goods sold = 720-276 = 444

  • So answer is b) $444